# reference.bib

### v2

@@ Acknowledgements
[…]
Appendix A Summary of results
@@
A.1 Two Stream example
Figure 7: Simulation of (2.1) with under-parametrized $H$ obtained from (2.6) using (KL) with far initialization using GD with linesearch. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
@@ A.1 Two Stream example
Figure 9: Simulation of (2.1) with under-parametrized $H$ obtained from (2.6) using (KL) with near initialization using GD with linesearch. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
@@
Figure 11: Simulation of (2.1) with under-parametrized $H$ obtained from (2.6) using (KL) with local initialization using GD with linesearch. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
@@
Figure 13: Simulation of (2.1) with under-parametrized $H$ obtained from (2.6) using (EE) with far initialization using GD with linesearch. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
@@
Figure 14: Simulation of (2.1) with under-parametrized $H$ obtained from (2.6) using (EE) with far initialization using GD with constant stepsize. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(T,x)$, $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$ and, convergence of objective.
@@
Figure 15: Simulation of (2.1) with under-parametrized $H$ obtained from (2.6) using (EE) with near initialization using GD with linesearch. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
@@
Figure 16: Simulation of (2.1) using (EE) with under-parametrized $H$ obtained from (2.6) with near initialization using GD with constant stepsize. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
@@
Figure 17: Simulation of (2.1) with under-parametrized $H$ obtained from (2.6) using (EE) with local initialization using GD with linesearch. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
@@
Figure 18: Simulation of (2.1) using (EE) with under-parametrized $H$ obtained from (2.6) with local initialization using GD with constant stepsize. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
@@
Figure 19: Simulation of (2.1) with over-parametrized $H$ obtained from (2.6) using (KL) with far initialization using GD with linesearch. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
@@
Figure 21: Simulation of (2.1) with over-parametrized $H$ obtained from (2.6) using (KL) with near initialization using GD with linesearch. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
@@
Figure 23: Simulation of (2.1) with over-parametrized $H$ obtained from (2.6) using (KL) with local initialization using GD with linesearch. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
@@
Figure 25: Simulation of (2.1) with under-parametrized $H$ obtained from (2.6) using (EE) with far initialization using GD with linesearch. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
@@
Figure 26: Simulation of (2.1) with over-parametrized $H$ obtained from (2.6) using (EE) with far initialization using GD with constant stepsize. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
@@
Figure 27: Simulation of (2.1) with over-parametrized $H$ obtained from (2.6) using (EE) with near initialization using GD with linesearch. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
@@
Figure 28: Simulation of (2.1) using (EE) with over-parametrized $H$ obtained from (2.6) with near initialization using GD with constant stepsize. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
@@
Figure 29: Simulation of (2.1) with over-parametrized $H$ obtained from (2.6) using (EE) with local initialization using GD with linesearch. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
@@
Figure 30: Simulation of (2.1) using (EE) with over-parametrized $H$ obtained from (2.6) with local initialization using GD with constant stepsize. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
A.2 Bump-on-Tail example
@@
Figure 31: Simulation of (2.1) with under-parametrized $H$ obtained from (2.6) using (KL) with far initialization using GD with linesearch. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).

### v3

@@ Acknowledgements
Appendix A Comparison between $L^{2}$ and KL
Another possible function is to use the $L^{2}$ disparity:
$\mathcal{J}(f[H])=\frac{1}{2}\|f_{T}[H]-f_{\text{eq}}\|_{L^{2}(x,v)}^{2}=\frac{1}{2}\int_{-L_{v}}^{L_{v}}(f[H](T,x,v)-f_{\text{eq}}(v))^{2}\mathrm{d}x\mathrm{d}v$ ($L^{2}$)
$\mathcal{J}(f[H])=\frac{1}{2}\int_{0}^{T}\int_{-L_{v}}^{L_{v}}(f[H](t,x,v)-f_{\text{eq}}(v))^{2}\mathrm{d}x\mathrm{d}v\mathrm{d}t\,.$ ($L^{2}$T)
In our simulation, we find the numerical behavior of these objective functions are rather similar to those of KL ((KL) and (KLT)). We provide numerical evidence in this section.
Figure 9: Landscape of Two Stream instability for $b_{1}$ for (4.1)(top) and Bump-on-Tail instability for $a_{1}$ for (4.3)(bottom) on the domain $[0.07,0.07]$ of the objectives (KL)(left), ($L^{2}$)(center-left), (KLT)(center-right) and ($L^{2}$T)(right).
In Figure 9, we show the one-dimensional case for both examples. For the Two Stream case, both landscapes exhibit no apparent differences, for both integrated-in-time case, and the disparity computed from the last frame. The situation for the Bump-on-Tail case is similar, with landscape for ($L^{2}$) agreeing with that for (KL) and ($L^{2}$T) agreeing with (KLT). We should note, however, ($L^{2}$) seems to show deeper local minima, making the computation more difficult to find the global optimal solution.
Appendix B Summary of results
@@
B.1 Two Stream example
Figure 10: Simulation of (2.1) with under-parametrized $H$ obtained from (2.6) using (KL) with far initialization using GD with line-search. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
@@ B.1 Two Stream example
Figure 12: Simulation of (2.1) with under-parametrized $H$ obtained from (2.6) using (KL) with near initialization using GD with line-search. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
@@
Figure 14: Simulation of (2.1) with under-parametrized $H$ obtained from (2.6) using (KL) with local initialization using GD with line-search. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
@@
Figure 16: Simulation of (2.1) with under-parametrized $H$ obtained from (2.6) using ($E$) with far initialization using GD with line-search. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
@@
Figure 17: Simulation of (2.1) with under-parametrized $H$ obtained from (2.6) using ($E$) with far initialization using GD with constant stepsize. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(T,x)$, $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$ and, convergence of objective.
@@
Figure 18: Simulation of (2.1) with under-parametrized $H$ obtained from (2.6) using ($E$) with near initialization using GD with line-search. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
@@
Figure 19: Simulation of (2.1) using ($E$) with under-parametrized $H$ obtained from (2.6) with near initialization using GD with constant stepsize. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
@@
Figure 20: Simulation of (2.1) with under-parametrized $H$ obtained from (2.6) using ($E$) with local initialization using GD with line-search. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
@@
Figure 21: Simulation of (2.1) using ($E$) with under-parametrized $H$ obtained from (2.6) with local initialization using GD with constant stepsize. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
@@
Figure 22: Simulation of (2.1) with under-parametrized $H$ obtained from (2.6) using (EET) with far initialization using GD with line-search. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
Figure 23: Simulation of (2.1) with under-parametrized $H$ obtained from (2.6) using (EET) with far initialization using GD with constant stepsize. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(T,x)$, $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$ and, convergence of objective.
Figure 24: Simulation of (2.1) with under-parametrized $H$ obtained from (2.6) using (EET) with near initialization using GD with line-search. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
Figure 25: Simulation of (2.1) using (EET) with under-parametrized $H$ obtained from (2.6) with near initialization using GD with constant stepsize. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
Figure 26: Simulation of (2.1) with under-parametrized $H$ obtained from (2.6) using (EET) with local initialization using GD with line-search. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
Figure 27: Simulation of (2.1) using (EET) with under-parametrized $H$ obtained from (2.6) with local initialization using GD with constant stepsize. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
Figure 28: Simulation of (2.1) with over-parametrized $H$ obtained from (2.6) using (KL) with far initialization using GD with line-search. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
@@
Figure 30: Simulation of (2.1) with over-parametrized $H$ obtained from (2.6) using (KL) with near initialization using GD with line-search. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
@@
Figure 32: Simulation of (2.1) with over-parametrized $H$ obtained from (2.6) using (KL) with local initialization using GD with line-search. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
@@
Figure 34: Simulation of (2.1) with under-parametrized $H$ obtained from (2.6) using ($E$) with far initialization using GD with line-search. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
@@
Figure 35: Simulation of (2.1) with over-parametrized $H$ obtained from (2.6) using ($E$) with far initialization using GD with constant stepsize. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
@@
Figure 36: Simulation of (2.1) with over-parametrized $H$ obtained from (2.6) using ($E$) with near initialization using GD with line-search. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
@@
Figure 37: Simulation of (2.1) using ($E$) with over-parametrized $H$ obtained from (2.6) with near initialization using GD with constant stepsize. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
@@
Figure 38: Simulation of (2.1) with over-parametrized $H$ obtained from (2.6) using ($E$) with local initialization using GD with line-search. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
@@
Figure 39: Simulation of (2.1) using ($E$) with over-parametrized $H$ obtained from (2.6) with local initialization using GD with constant stepsize. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
@@
Figure 40: Simulation of (2.1) with under-parametrized $H$ obtained from (2.6) using (EET) with far initialization using GD with line-search. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
Figure 41: Simulation of (2.1) with over-parametrized $H$ obtained from (2.6) using (EET) with far initialization using GD with constant stepsize. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
Figure 42: Simulation of (2.1) with over-parametrized $H$ obtained from (2.6) using (EET) with near initialization using GD with line-search. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
Figure 43: Simulation of (2.1) using (EET) with over-parametrized $H$ obtained from (2.6) with near initialization using GD with constant stepsize. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
Figure 44: Simulation of (2.1) with over-parametrized $H$ obtained from (2.6) using (EET) with local initialization using GD with line-search. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
Figure 45: Simulation of (2.1) using (EET) with over-parametrized $H$ obtained from (2.6) with local initialization using GD with constant stepsize. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
B.2 Bump-on-Tail example
Figure 46: Simulation of (2.1) with under-parametrized $H$ obtained from (2.6) using (KL) with far initialization using GD with line-search. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).